# RefinedFields: Radiance Fields Refinement for Unconstrained Scenes

### v3

@@ 3.2 Scene Refining
Given a fitted scene representation $\mathbf{P}_{\gamma}$, this stage consists of learning this fitted implicit representation and proposing a new refined representation. Formally, this stage consists of projecting our K-Planes $\mathbf{P}_{\gamma}$ on the set $\mathbb{Q}$ of K-Planes inferable by a low-rank fine-tuning of the pre-trained model:
@@
As K-Planes feature channels show similar structure to images (Appendix C), this projection pushes the K-Planes to be even more similar in structure to real images, more particularly to orthogonal projections of the scene on the planes. This leads to a better initialization for the K-Planes optimization, as proven by our experiments (Figures 3 and 2).
To provide scene refining with a rich prior, we employ a large-scale pre-trained latent diffusion model, as these networks exhibit great performances as priors for downstream tasks, and share similar properties to our planar representation, both in terms of shape and distribution (Appendix C). More particularly, we adopt Stable Diffusion (Rombach et al., 2022, SD) for its proven performances for downstream 3D (Liu et al., 2023) and 2D (Wang et al., 2023) tasks. Thus, we integrate the U-Net $\mathbf{SD}_{\phi}$ and the decoder $\mathbf{D}_{\omega}$ into our pipeline, and treat the K-Planes as $3C$-channel $N\times N$ images. We also replace the last layer of the decoder $\mathbf{D}_{\omega}$ with a randomly initialized convolutional layer (with no bias), to take into account the shape of the K-Planes. We then fine-tune the pre-trained model using the fitted K-Planes $\mathbf{P}_{\gamma}$, and infer refined K-Planes $\mathbf{P}_{\varepsilon}=\mathbf{D}_{w}(\mathbf{SD}_{\phi}(x))$ where $x\sim\mathcal{N}(0,I)$. Note that this is different from the multi-step generation process of diffusion model inference, as we only apply the inference at the last time-step of our diffusion model. This is key as our goal here is not to learn distributions over scenes and sample them for generation, but to adapt the pre-trained network and leverage the information already learned within its weights to infer K-Planes closest to representing the scene at hand.
@@
To achieve the fine-tuning of our pre-trained network, a significant challenge presents itself: due to the sheer size of Stable Diffusion, it would be exceptionally costly to fine-tune all of its trainable parameters. Moreover, as we only want to modulate priors embedded into the pre-trained network, we look for an alternative to doing full fine-tuning. To circumvent these constraints, we adopt Low-Rank Adaptation (Hu et al., 2022, LoRA), a simple yet effective parameter-efficient fine-tuning method that has proven great transfer capabilities across modalities and tasks (Fan et al., 2023; Lee et al., 2023; Zeng & Lee, 2023). LoRA’s relatively minimal design works directly over weight tensors, which means that it can be seamlessly applied to most model architectures. Furthermore, LoRA does not add any additional cost at inference, thanks to its structural re-parameterization design. To achieve this, Hu et al. (2022) inject trainable low-rank decomposition matrices into each layer of a frozen pre-trained model. Let $\mathbf{W}_{0}$, $\mathbf{b}_{0}$ be the frozen pre-trained weights and biases, and $x$ be the input. Fine-tuning a frozen linear layer $f(x)=\mathbf{W}_{0}x+\mathbf{b}_{0}$ comes down to learning the low-rank decomposition weights $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$:
$f(x)=(\mathbf{W}_{0}+\Delta\mathbf{W})x+\mathbf{b}_{0}$ (6)
Thus, to implement scene refining, we fine-tune the LoRA parameters $\phi$ modulating the pre-trained U-Net, as well as the decoder’s parameters $\omega$, on the fitted scene $\mathbf{P}_{\gamma}$. Subsequently, we query the U-Net with Gaussian noise $x$, decode its intermediary output latent $z$ with $\mathbf{D}_{\omega}$, and infer $\mathbf{p}_{\varepsilon}$ that is reshaped into a new refined scene $\mathbf{P}_{\varepsilon}$. Finally, $\mathbf{P}_{\varepsilon}$ is proposed to scene fitting as an improved initialization to be optimized. For an in-depth inspection of the feature planes, we highly encourage the reader to refer to Appendix C.
@@ 3.3 Training
[…]
@@
where $x\sim\mathcal{N}(0,I)$ is fixed during scene refining, $\mathbf{SD}_{\phi}$ is the frozen Stable Diffusion model modulated by LoRA with trainable parameters $\phi$, and $\mathbf{D}_{w}$ is the latent K-Planes decoder. After this optimization, $\mathbf{P}_{\gamma}$ is reassigned as $\mathbf{D}_{w}(\mathbf{SD}_{\phi}(x))$ and passed to scene fitting. Note that, thanks to the alternating nature of our training and the absence of bias in the decoder’s convolutional layers, this optimization does not reach full convergence. This is key because over-fitting the generative model to generate exactly $\mathbf{P}_{\gamma}$ leads to resuming scene fitting from exactly the same point.
@@ 4 Experiments
Table 2: Quantitative results. Results on three real-world datasets from Phototourism (Jin et al., 2020). The bold and underlined entries respectively indicate the best and second-best results. Our method shows notable improvements compared to K-Planes in NVS-W.
Brandenburg Gate Sacré Coeur Trevi Fountain PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) NeRF (Mildenhall et al., 2020) 18.90 0.8159 15.60 0.7155 16.14 0.6007 NeRF-W (Martin-Brualla et al., 2021) 24.17 0.8905 19.20 0.8076 18.97 0.6984 K-Planes (Fridovich-Keil et al., 2023) 25.49 0.8785 20.61 0.7735 22.67 0.7139 K-Planes-SS (Fridovich-Keil et al., 2023) 24.48 0.8629 19.86 0.7419 21.30 0.6627 RefinedFields-noFinetuning (ours) 25.39 0.8834 21.41 0.8059 22.54 0.7324 RefinedFields-noPrior (ours) 25.42 0.8822 21.17 0.7978 22.16 0.7251 RefinedFields (ours) 26.64 0.8869 22.26 0.8176 23.42 0.7379
We start by assessing RefinedFields via an experiment on a synthetic example. We then evaluate RefinedFields on real-world Phototourism (Jin et al., 2020) scenes, where we showcase the improvements our method exhibits relative to our K-Planes base representation. Quantitative results can be found in Table 2, where we report for each experiment the Peak Signal-to-Noise Ratio (PSNR) for pixel-level similarity and the Structural Similarity Index Measure (SSIM) for structural-level similarity. RefinedFields demonstrates superior performance compared to K-Planes on the task of novel view synthesis in-the-wild. For a further look, experimental details including hyperparameters and more dataset details can be found in Appendix B. Quantitative results on addtional synthetic scenes can be found in Appendix D. Additional real-world qualitative results on Phototourism (Jin et al., 2020) are also available in Appendix D.
@@ 4.1 Datasets
We evaluate our method similarly to prior work (Martin-Brualla et al., 2021; Chen et al., 2022b; Fridovich-Keil et al., 2023) in novel view synthesis in-the-wild, by adopting the same three scenes of cultural monuments from the Phototourism dataset (Jin et al., 2020): Brandenburg Gate, Sacré Coeur, and Trevi Fountain. Additional dataset details can be found in Appendix B.
4.2 Implementation Details
For a fair comparison, we take similar experimental settings in scene fitting to Fridovich-Keil et al. (2023). However, due to the nature of our scene refining pipeline, we limit the implementation in our case to a single-scale K-Planes of $512\times 512$ resolution, in contrast to the multi-scale approach taken by Fridovich-Keil et al. (2023) where $N\in\{64,128,256,512\}$. The number of channels in each plane remains the same ($C=32$). Also note that throughout all the experiments, we consider the hybrid implementation of K-Planes, where plane features are decoded into colors and densities by a small MLP. As for the scene refining pipeline, we apply no modification to the U-Net in Stable Diffusion. Yet, we replace the last layer of the decoder $D_{\omega}$ with a new convolutional layer (without bias) to account for the shape of the K-Planes. Hence, the dimensions we work with in scene refining (Figure 2) are: $N=512$, $C=32$, $n=64$, and $c=4$. For an in-depth look at our hyperparameter settings, we refer the reader to Appendix B.
4.3 Evaluations
Baselines.
We compare our results against NeRF-W (Martin-Brualla et al., 2021) (results from the public implementation (Aoi, 2022)) and K-Planes (Fridovich-Keil et al., 2023). Note that, to assess the added value of our refining pipeline with respect to our base representation (Section 4.2), we also include a single-scale ablation of K-Planes with $N=512$ (dubbed K-Planes-SS).
Comparisons.
We start by testing our method against K-Planes on a toy example consisting of the Lego scene from the NeRF synthetic dataset. Here, we train both methods on half of the training set. As illustrated in Figure 3 our method refines K-Planes and exhibits better quantitative and qualitative results thanks to our scene refining pipeline.
We then apply our method on in-the-wild scenes. In this case, RefinedFields outperforms not only K-Planes-SS but also K-Planes and NeRF-W on the task of NVS-W (Table 2). This particularly highlights the value of scene refining. Figures 1 and 4 show qualitative comparisons of RefinedFields with K-Planes, showing the visual improvements brought by our refining pipeline, which brings finer details to monuments in the Phototourism scenes. Further qualitative results can be found in Appendix D.
4.4 Ablations
To justify our choices and explore further, we compare our in-the-wild results (Table 2) to results from two main ablations of our method. RefinedFields-noFinetuning is a variation of our method without LoRA fine-tuning. Here, we consider the same exact pipeline (frozen U-Net, same decoder configuration), except that we don’t modulate the weights of the frozen U-Net with LoRA. This means that the prior is kept intact and no fine-tuning is done. This is to assess the role that LoRA finetuning of the pre-trained prior plays in our pipeline. RefinedFields-noPrior ablates the prior of Stable Diffusion by randomly re-initializing all U-Net weights while leaving all other elements of the scene-refining pipeline intact. This ablation is done to evaluate the importance of the prior, and to verify that the observed refinements are not entirely resulting from alternate training. Note that ablating the entire scene refining pipeline leads back to the K-Planes-SS setting. As illustrated in Table 2, we consistently obtain worse results during the ablation study as compared to our full model, thus demonstrating the value of the pre-trained prior and of LoRA finetuning. This proves the importance of extending scene learning beyond closed-world settings, as our optimization conditioning comes from a large image prior that a reasonably sized training set cannot fully capture, especially in-the-wild.
4.5 Training Time Analysis
As presented, RefinedFields utilizes an alternating training procedure and a pre-trained prior to refine scene representations, leading to richer details in rendered images. While our method demonstrates interesting results, this however comes with a training time increase as compared to our base representation, as our alternating training leads to the repeated fine-tuning of both scene fitting and scene refining pipelines. In addition, in order for this training to converge well, we apply a relatively small learning rate ($10^{-4}$) to the LoRA parameters. While we leave training time optimizations for future work, our method still requires significantly less training time compared to its concurrent works in novel view synthesis in-the-wild (Table 3).
Table 3: Training times. Comparison of training time among methods tackling novel view synthesis in-the-wild. Timings are based on a single NVIDIA V100 GPU. Our method improves upon K-Planes while still requiring significantly less training time than other works.
Methods trained in a half-resolution setting.
Training Time ($\downarrow$)
NeRF-W (Martin-Brualla et al., 2021) 400 hrs
[…]
K-Planes (Fridovich-Keil et al., 2023) 50 mins
RefinedFields (ours) 150 hrs
5 Conclusion
In this paper, we introduce RefinedFields, a method that refines K-Planes representations by incorporating a pre-trained prior and alternate training . Extensive experiments show that RefinedFields exhibits notable improvements on real-world monument details. In concluding this study, several avenues of future work emerge as we consider this work to be a first stepping-stone in conditioning closed-world in-the-wild scene representations with extrinsic signals. This includes the exploration of approaches to achieve this conditioning other than optimization guidance, the application of scene refining on other representations, and the extension of appearance modeling beyond learnable embeddings.
Impact Statement
This paper presents work that enhances the construction of high-quality neural representations. As such, the risks associated with our work parallel those of other neural rendering papers. This includes but is not limited to privacy and security concerns, as our method is trained on a dataset of publicly captured images, where privacy-sensitive information (e.g. human faces, license plate numbers) could be present. Hence, similarly to other neural rendering approaches, there is a risk that such data could end up in the trained model if the employed datasets are not properly filtered before use. Furthermore, as our work utilizes Stable Diffusion as prior, it inherits any problematic biases and limitations this model may have (Carlini et al., 2023; Luccioni et al., 2023).
Acknowledgements
Appendix A Related Work Overview
Due to space constraints, this section presents the referenced version of Table 1 in the main paper.
Table 4: Related work overview. RefinedFields leverages unconstrained tourist photo collections (Jin et al., 2020) and a pre-trained prior (Rombach et al., 2022) to refine K-Planes, our underlying scene representation.
NeRF (Mildenhall et al., 2020) NeRF NVS
K-Planes (Fridovich-Keil et al., 2023) K-Planes
NFD (Shue et al., 2023) Tri-Planes OG
3DGen (Gupta et al., 2023) Tri-Planes
Latent-NeRF (Metzer et al., 2023) NeRF
DreamFusion (Poole et al., 2023) NeRF
PixelNeRF (Yu et al., 2021) NeRF NVS
NeRDi (Deng et al., 2023) NeRF
RealFusion (Melas-Kyriazi et al., 2023) NeRF
DiffusioNeRF (Wynn & Turmukhambetov, 2023) NeRF
NerfDiff (Gu et al., 2023) Tri-Planes
3DiM (Watson et al., 2023)
Zero-1-to-3 (Liu et al., 2023)
NeRF-W (Martin-Brualla et al., 2021) NeRF NVS-W
Ha-NeRF (Chen et al., 2022b) NeRF
CR-NeRF (Yang et al., 2023) NeRF
RefinedFields (ours) K-Planes
Appendix B Experimental Details
B.1 Datasets
Synthetic dataset.
In-the-wild dataset.
B.2 Frameworks
We make use of multiple frameworks to implement our method. Our Python source code (tested on version 3.7.16), based on PyTorch (Paszke et al., 2019) (tested on version 1.13.1) and CUDA (tested on version 11.6), will be publicly available as open source. We also utilize Diffusers (von Platen et al., 2022) and Stable Diffusion (tested on version 1-5, main revision). K-Planes also adopt the tinycudann framework (Müller, 2021). We run all experiments on a single NVIDIA V100 GPU.
B.3 Hyperparameters
Appendix C Feature Planes Inspection
Appendix D Supplementary Results
Quantitative results.
We reproduce the results on K-Planes and K-Planes-SS. In-the-wild results on NeRF-W come from a public implementation of the paper (Aoi, 2022) and are extracted from the K-Planes paper. Remaining results in Tables 5 and 2 are extracted from their corresponding papers.
Qualitative results.
We present below additional qualitative results for in-the-wild scenes (Figures 8, 9 and 10) and synthetic renderings Figures 11 and 12. Please refer to our project page for animations of our renderings.
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes

### v4

@@ 3.2 Scene Refining
This section presents the core of our method, which consists of proposing better optimization initializations for the scene fitting stage. Given a fitted scene representation $\mathbf{P}_{\gamma}$, this stage consists of learning this fitted implicit representation and proposing a new refined representation $\mathbf{P}_{\varepsilon}$. Formally, this stage consists of projecting our K-Planes $\mathbf{P}_{\gamma}$ on the set $\mathbb{Q}$ of K-Planes inferable by a low-rank fine-tuning of the pre-trained model:
@@
As K-Planes feature channels show similar structure to images (Appendix B), this projection pushes the K-Planes to be even more similar in structure to real images, more particularly to orthogonal projections of the scene on the planes. Figures 5, 6 and 7 illustrate a comparison between feature planes at the end of our optimization and those of standard K-Planes. RefinedFields leads to feature planes exhibiting sharper details, which ultimately lead to more refined details in scene renderings, as proven by our experiments (Figures 3 and 3).
To provide scene refining with a rich prior, we employ a large-scale pre-trained latent diffusion model, as these networks exhibit great performances as priors for downstream tasks, and share similar properties to our planar representation, both in terms of shape and distribution (Appendix B). More particularly, we adopt Stable Diffusion (Rombach et al., 2022, SD) for its proven performances for downstream 3D (Liu et al., 2023) and 2D (Wang et al., 2023a) tasks. Thus, we integrate the U-Net $\mathbf{SD}_{\phi}$ and the decoder $\mathbf{D}_{\omega}$ into our pipeline, and treat the K-Planes as $3C$-channel $N\times N$ images. We also replace the last layer of the decoder $\mathbf{D}_{\omega}$ with a randomly initialized convolutional layer (with no bias), to take into account the shape of the K-Planes. We then fine-tune the pre-trained model using the fitted K-Planes $\mathbf{P}_{\gamma}$, and infer refined K-Planes $\mathbf{P}_{\varepsilon}=\mathbf{D}_{w}(\mathbf{SD}_{\phi}(x))$ where $x\sim\mathcal{N}(0,I)$ is sampled once at the beginning of the training. Note that this is different from the multi-step generation process of diffusion model inference, as we only apply the inference at the last time-step of our diffusion model. This is key as our goal here is not to learn distributions over scenes and sample them for generation, but to adapt the pre-trained network and leverage the information already learned within its weights to infer K-Planes closest to representing the scene at hand.
To achieve the fine-tuning of our pre-trained network, a significant challenge presents itself: due to the sheer size of Stable Diffusion, it would be too costly to fine-tune all of its trainable parameters. Moreover, as we only want to modulate priors embedded into the pre-trained network, we look for an alternative to doing full fine-tuning. To circumvent these constraints, we adopt Low-Rank Adaptation (Hu et al., 2022, LoRA), a simple yet effective parameter-efficient fine-tuning method that has proven great transfer capabilities across modalities and tasks (Fan et al., 2023; Lee et al., 2023; Zeng & Lee, 2023). LoRA’s relatively minimal design works directly over weight tensors, which means that it can be seamlessly applied to most model architectures. Furthermore, LoRA does not add any additional cost at inference, thanks to its structural re-parameterization design. To achieve this, Hu et al. (2022) inject trainable low-rank decomposition matrices into each layer of a frozen pre-trained model. Let $\mathbf{W}_{0}$, $\mathbf{b}_{0}$ be the frozen pre-trained weights and biases, and $x$ be the input. Fine-tuning a frozen linear layer $f(x)=\mathbf{W}_{0}x+\mathbf{b}_{0}$ comes down to learning the low-rank decomposition weights $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$:
$f(x)=(\mathbf{W}_{0}+\Delta\mathbf{W})x+\mathbf{b}_{0}$ (6)
Thus, to implement scene refining, we fine-tune the LoRA parameters $\phi$ modulating the pre-trained U-Net, as well as the decoder’s parameters $\omega$, on the fitted scene $\mathbf{P}_{\gamma}$. Subsequently, we query the U-Net with Gaussian noise $x$, decode its intermediary output latent $z$ with $\mathbf{D}_{\omega}$, and infer $\mathbf{p}_{\varepsilon}$ that is reshaped into a new refined scene $\mathbf{P}_{\varepsilon}$. Finally, $\mathbf{P}_{\varepsilon}$ is proposed to scene fitting as an improved initialization to be optimized. For an in-depth inspection of the feature planes, we refer the reader to Appendix B.
@@ 3.3 Training
$\mathcal{L}_{\mathrm{TV}}(\mathbf{P})=\frac{1}{|C|N^{2}}\sum_{c,i,j}(\lVert\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i-1,j}\rVert_{2}^{2}+\lVert\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i,j-1}\rVert_{2}^{2})~{}.$ (7)
@@
where $x\sim\mathcal{N}(0,I)$ is fixed during scene refining, $\mathbf{SD}_{\phi}$ is the frozen Stable Diffusion model modulated by LoRA with trainable parameters $\phi$, and $\mathbf{D}_{w}$ is the latent K-Planes decoder. After this optimization, $\mathbf{P}_{\gamma}$ is reassigned as $\mathbf{D}_{w}(\mathbf{SD}_{\phi}(x))$ and passed to scene fitting. Note that, thanks to the alternating nature of our training and the absence of bias in the decoder’s convolutional layers, this optimization does not overfit the model to produce exactly $\mathbf{P}_{\gamma}$, which is key as it would lead to resuming scene fitting from exactly the same point.
@@
Figure 3: Case study. Qualitative results on the Lego scene from the NeRF synthetic dataset (Mildenhall et al., 2020) showcasing the optimization progression on RefinedFields, and a comparison with the ground truth and K-Planes. The training set is constrained to 50% of its initial size for both RefinedFields and K-Planes. RefinedFields refines the K-Planes representation enabling the proper reconstruction of details in the scene. At the end of optimization, the Mean Squared Error (MSE) for RefinedFields is $3.46\times 10^{-4}$, while the one for K-Planes is $4.36\times 10^{-4}$.
@@ 4 Experiments
We start by assessing RefinedFields via an experiment on a case study. We then evaluate RefinedFields on synthetic scenes (Mildenhall et al., 2020) and real-world Phototourism (Jin et al., 2020) scenes, where we showcase the improvements our method exhibits relative to our K-Planes base representation. Quantitative results can be found in Tables 2 and 3, where we report for each experiment the Peak Signal-to-Noise Ratio (PSNR) for pixel-level similarity, the Structural Similarity Index Measure (SSIM) for structural-level similarity, and the Learned Perceptual Image Patch Similarity (Zhang et al., 2018, LPIPS) for perceptual similarity. RefinedFields demonstrates an improved performance compared to K-Planes on the task of novel view synthesis. For a further look, experimental details including hyperparameters and more dataset details can be found in Appendix A. Additional qualitative results on synthetic and real-world scenes are available in Appendix C.
@@ 4.1 Datasets
We evaluate our method similarly to prior work (Mildenhall et al., 2020; Martin-Brualla et al., 2021; Fridovich-Keil et al., 2023) in novel view synthesis, by adopting the Real Synthetic $\mathit{360^{\circ}}$ dataset (Mildenhall et al., 2020) for synthetic scenes and the same three scenes of cultural monuments from the Phototourism dataset (Jin et al., 2020) for real-world scenes: Brandenburg Gate, Sacré Coeur, and Trevi Fountain. Additional dataset details can be found in Section A.1.
@@
4.2 Implementation Details
For a fair comparison, we take similar experimental settings in scene fitting to Fridovich-Keil et al. (2023). However, due to the nature of our scene refining pipeline, we limit the implementation in our case to a single-scale K-Planes of $512\times 512$ resolution, in contrast to the multi-scale approach taken by Fridovich-Keil et al. (2023) where $N\in\{64,128,256,512\}$. The number of channels in each plane remains the same ($C=32$). Moreover, throughout all the experiments, we consider the hybrid implementation of K-Planes, where plane features are decoded into colors and densities by a small MLP. As for the scene refining pipeline, we apply no modification to the U-Net in Stable Diffusion. Yet, we replace the last layer of the decoder $D_{\omega}$ with a new convolutional layer (without bias) to account for the shape of the K-Planes. Hence, the dimensions used in scene refining (Figure 2) are: $N=512$, $C=32$, $n=64$, and $c=4$. For an in-depth look at our frameworks and hyperparameter settings, we refer the reader to Sections A.2 and A.3.
4.3 Evaluations
Baselines.
For both synthetic and in-the-wild scenes, we primarily compare our method to our K-Planes baseline, highlighting the improvements it brings to planar scene representations. Note that, for a fair comparison, and to assess the added value of our refining pipeline with respect to our base representation (Section 4.2), we also include a single-scale ablation of K-Planes with $N=512$ (dubbed K-Planes-SS). In order to provide a more comprehensive perspective, we also illustrate the NVS performances of other recent works that do not employ planar scene representations for synthetic (Mildenhall et al., 2020; Chen et al., 2022a; Fridovich-Keil et al., 2022; Müller et al., 2022) and in-the-wild (Mildenhall et al., 2020; Martin-Brualla et al., 2021; Xu et al., 2024; Kulhanek et al., 2024) scenes.
Comparisons.
We start by testing our method against K-Planes on a case study consisting of the Lego scene from the NeRF synthetic dataset. Here, we train both methods on half of the training set as to deliberately produce a lower-quality fitted scene. As illustrated in Figure 3 our method refines K-Planes and exhibits better quantitative and qualitative results thanks to our scene refining pipeline.
We then apply our method to learn synthetic and in-the-wild scenes. In this case, RefinedFields improves upon not only K-Planes-SS but also K-Planes on the task of NVS (Tables 2 and 3). As a result, it enhances the performance of planar scene representations and brings them closer to recent works, which highlights the value of scene refining. Particularly, for synthetic scenes, RefinedFields improves upon K-Planes and sometimes even outperforms other state-of-the-art methods. For in-the-wild scenes, RefinedFields also improves K-Planes performances in a notable way. However, recent state-of-the-art methods based on Gaussian Splatting architectures (Kerbl et al., 2023) continue to outperform the refined planar representations in terms of overall quality. Figures 1 and 4 show qualitative comparisons of RefinedFields with K-Planes, showing the visual improvements brought by our refining pipeline, which brings finer details to monuments in the Phototourism scenes. Further qualitative results on synthetic and in-the-wild scenes can be found in Appendix C. We also present an inspection of K-Planes features learned by our method and K-Planes in Appendix B.
It is important to note that we do not compare our NVS metrics with other recent works (Chen et al., 2022b; Yang et al., 2023; Zhang et al., 2025). This is because these methods are trained and evaluated on downscaled versions of the Phototourism dataset, which reduces the prominence of fine details in the ground truth for monument structures. As their NVS metrics are computed relative to these ground truth, their results are not comparable to ours. Although other works do compare NVS metrics of various methods across different resolutions, we refrain from this practice as it is not an accurate comparison.
Table 3: Quantitative results. Results on three real-world datasets from Phototourism (Jin et al., 2020). Our method shows notable improvements compared to K-Planes on the task of NVS-W.
Results from public implementation (Aoi, 2022) reproduced by Fridovich-Keil et al. (2023).
Brandenburg Gate Sacré Coeur Trevi Fountain PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) NeRF 18.90 0.8159 0.231 15.60 0.7155 0.291 16.14 0.6007 0.366 NeRF-W 21.32 19.17 18.61 Splatfacto-W 26.87 0.9320 0.124 22.53 0.8760 0.158 22.66 0.7690 0.224 WildGaussians 27.77 0.9270 0.133 22.56 0.8590 0.177 23.63 0.7660 0.228 K-Planes 25.49 0.8785 0.224 20.61 0.7735 0.265 22.67 0.7139 0.317 K-Planes-SS 24.48 0.8629 0.242 19.86 0.7419 0.312 21.30 0.6627 0.355 RefinedFields-noFinetuning 25.39 0.8834 0.206 21.41 0.8059 0.239 22.54 0.7324 0.291 RefinedFields-noPrior 25.42 0.8822 0.214 21.17 0.7978 0.248 22.16 0.7251 0.291 RefinedFields (ours) 26.64 0.8869 0.206 22.26 0.8176 0.228 23.42 0.7379 0.284
As presented, RefinedFields utilizes an alternating training procedure and a pre-trained prior to refine scene representations, leading to richer details in rendered images. While our method demonstrates promising results, this however comes with a training time increase as compared to our base representation, as our K-Planes feature projection via alternating training leads to the repeated fine-tuning of both scene fitting and scene refining pipelines, which takes overall around 80 hours on a single NVIDIA A100 GPU. We leave the optimization of training time for future work.
4.4 Ablations
To justify our choices and explore further, we compare our in-the-wild results (Table 3) to results from two main ablations of our method. RefinedFields-noFinetuning is a variation of our method without LoRA fine-tuning. Here, we consider the same exact pipeline (frozen U-Net, same decoder configuration), except that we don’t modulate the weights of the frozen U-Net with LoRA. This means that the prior is kept intact and no fine-tuning is done. This is to assess the role that LoRA finetuning of the pre-trained prior plays in our pipeline. Note that the decoder is still fine-tuned in this ablation, as additionally freezing it would lead to a random constant optimization initialization $\mathbf{P}_{\varepsilon}$, which would defeat the purpose of the refining stage. RefinedFields-noPrior ablates the prior of Stable Diffusion by randomly re-initializing all U-Net weights while leaving all other elements of the scene-refining pipeline intact. This ablation is done to evaluate the importance of the prior, and to verify that the observed refinements are not entirely resulting from alternate training. Note that ablating the entire scene refining pipeline leads back to the K-Planes-SS setting. As illustrated in Table 3, we consistently obtain worse results during the ablation study as compared to our full model, thus demonstrating the value of the pre-trained prior and of LoRA finetuning. This proves the importance of extending scene learning beyond closed-world settings, as our K-Planes projection is done via an optimization conditioning coming from a large image prior that a reasonably sized training set cannot fully capture, especially in-the-wild.
5 Conclusion
In this paper, we introduce RefinedFields, a method that refines K-Planes representations by using a pre-trained prior and an alternate training procedure. Extensive experiments show that RefinedFields exhibits notable improvements on the task of novel view synthesis compared to its K-Planes baseline. In concluding this study, several avenues of future work emerge as we consider this work to be a first stepping-stone in improving planar scene representations via conditioning with extrinsic signals. This includes the exploration of approaches to achieve this conditioning other than optimization guidance, and the application of scene refining on other representations.
Impact Statement
This paper presents work that enhances the construction of high-quality neural representations. As such, the risks associated with our work parallel those of other neural rendering papers. This includes but is not limited to privacy and security concerns, as our method is trained on a dataset of publicly captured images, where privacy-sensitive information (e.g. human faces, license plate numbers) could be present. Hence, similarly to other neural rendering approaches, there is a risk that such data could end up in the trained model if the employed datasets are not properly filtered before use. Furthermore, as our work utilizes Stable Diffusion as prior, it inherits any problematic biases and limitations this model may have.
Acknowledgments
Appendix A Experimental Details
A.1 Datasets
Synthetic dataset.
In-the-wild dataset.
A.2 Frameworks
We make use of multiple frameworks to implement our method. Our Python source code (tested on version 3.7.16), based on PyTorch (Paszke et al., 2019) (tested on version 1.13.1) and CUDA (tested on version 11.6), is publicly available as open source. We also utilize Diffusers (von Platen et al., 2022) and Stable Diffusion (tested on version 1-5, main revision). K-Planes also adopt the tinycudann framework (Müller, 2021). We run all experiments on a single NVIDIA A100 GPU.
A.3 Hyperparameters
Appendix B Feature Planes Inspection
Appendix C Supplementary Results
We present below additional qualitative results for in-the-wild scenes (Figures 8, 9 and 10) and synthetic scenes (Figures 11 and 12).
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes
(a) RefinedFields (In-Process)
(b) RefinedFields (Final)
(c) K-Planes